# Language models scale reliably with over-training and on downstream tasks

### 13 Mar 2024 v1

@@ Abstract
Scaling laws are useful guides for developing language models, but there are still gaps between current scaling studies and how language models are ultimately trained and evaluated. For instance, scaling is usually studied in the compute-optimal training regime (i.e., “Chinchilla optimal” regime); however, in practice, models are often over-trained to reduce inference costs. Moreover, scaling laws mostly predict loss on next-token prediction, but ultimately models are compared based on downstream task performance. In this paper, we address both shortcomings. To do so, we create a testbed of 104 models with 0.011B to 6.9B parameters trained with various numbers of tokens on three data distributions. First, we investigate scaling in the over-trained regime. We fit scaling laws that extrapolate in both the number of model parameters and the ratio of training tokens to parameters. This enables us to predict the validation loss of a 1.4B parameter, 900B token run (i.e., 32$\times$ over-trained) and a 6.9B parameter, 138B token run—each from experiments that take 300$\times$ less compute. Second, we relate the perplexity of a language model to its downstream task performance via a power law. We use this law to predict top-1 error averaged over downstream tasks for the two aforementioned models using experiments that take 20$\times$ less compute. Our experiments are available at https://github.com/mlfoundations/scaling.
00footnotetext: Equal advising, ordered alphabetically. Correspondence to sy@cs.columbia.edu. ${}^{1}$olumbia University ${}^{2}$oyota Research Insitute ${}^{3}$ Austin ${}^{4}$pple ${}^{5}$niversity of Washington ${}^{6}$uelich Supercomputing Center, Research Center Juelich ${}^{7}$ ${}^{8}$tanford University ${}^{9}$el Aviv University ${}^{10}$ Munich ${}^{11}$ontextual AI
@@ 1 Introduction
Figure 1: Reliable scaling in the over-trained regime and for downstream error prediction. (left) We fit a scaling law for model validation loss, parameterized by (i) a token multiplier $M$, which is the ratio of training tokens $D$ to parameters $N$ (i.e., $M=D/N$) and (ii) the compute $C$ in FLOPs used to train a model, approximated by $C=6ND$. We extrapolate, in both $N$ and $M$, the validation performance of models requiring over 300$\times$ the training compute used to construct the scaling law. (right) We also fit a scaling law to predict average downstream top-1 error as a function of validation loss. We find that fitting scaling laws for downstream error benefits from using more expensive models when compared to fitting for loss prediction. We predict the average error over 17 downstream tasks for models trained with over 20$\times$ the compute.
@@
Training large language models is expensive. Moreover, training high-quality models requires a complex recipe of algorithmic techniques and training data. To reduce the cost of finding successful training recipes, researchers first evaluate ideas with small experiments and then extrapolate their efficacy to larger scales. With reliable extrapolation, it is possible to quickly iterate at small scale and still pick the method that will perform best for the final large training run. Indeed, this workflow has become commonplace for training state-of-the-art language models such as Chinchilla 70B [43], PaLM 540B [17], and GPT-4 [74].
@@
Despite their importance for model development, published scaling laws differ from the goals of training state-of-the-art models in important ways. For instance, scaling studies usually focus on the compute-optimal training regime (“Chinchilla optimality” [43]), while widely used models are now often over-trained to reduce inference costs. Another potential mismatch between scaling laws and eventual applications of the models is that most scaling laws quantify model performance by perplexity in next-token prediction instead of accuracy on widely used benchmark datasets. As a result, it is unclear whether following scaling laws leads to truly better models, or merely to models with lower perplexity in the compute-optimal training regime.
@@
In this paper, we address both topics: scaling in the over-trained regime and downstream performance prediction, with an extensive set of experiments.
@@
Motivated by the practice of training beyond compute-optimal, we first investigate whether scaling follows reliable trends in the over-trained regime. We find that for a set of model configurations with a constant ratio of training tokens to parameters, the models’ reducible loss $L^{\prime}$ [41, 43] follows consistent power laws ($L^{\prime}=\lambda\cdot C^{-\alpha_{C}}$) in the amount of training compute $C$. As one increases the ratio of tokens to parameters, corresponding to more over-training, the scaling exponent $\alpha_{C}$ remains about the same, while the scalar $\lambda$ changes. We consider the extent to which our observations are explainable and find a promising approach by reparameterizing forms for scaling laws.
@@
To establish if and when scaling is predictable in the over-trained regime, we experiment with a testbed of 104 models, trained from scratch on three different datasets: RedPajama [110], C4 [86, 25], and RefinedWeb [80]. We find that scaling laws fit on small models, trained closer to compute-optimal, can accurately predict the performance of larger models that undergo more over-training. Figure 1 (left) illustrates our main over-training result, where we invest $192.4e$ 19 FLOPs to extrapolate the C4 validation performance of a 1.4B parameter model trained on 900B tokens, which requires 300×300$\times$300 × more compute to train.
@@
In addition to over-training, we also investigate if scaling laws can predict the performance of a model on downstream tasks. We establish a power law relationship between language modeling perplexity and the average top-1 error on a suite of downstream tasks. While it can be difficult to predict the error on individual tasks, we find that aggregate performance can be accurately predicted from a model’s perplexity among models trained on the same training data. Figure 1 (right) presents our main downstream error prediction result, where we invest $202.7e$ 20 FLOPs to predict the average top-1 error over a set of downstream tasks to within 1 percentage point for a 6.9B compute-optimal model, which requires 20×20$\times$20 × more compute to train.
@@
To facilitate further research on reliable scaling, we provide all results of our experiments at https://github.com/mlfoundations/scaling.
2 Scaling and over-training
In this section, we describe empirical observations and their potential mathematical descriptions. First, we provide key definitions (Section 2.1). We next present a phenomenon wherein training a collection of models for increasing token multipliers—ratios of tokens to parameters—follows similar scaling trends (Section 2.2). We then show that these patterns are consistent with previously proposed power-law scaling when reparameterizing in terms of training compute and token multipliers (Section 2.3). Towards connecting loss scaling and downstream performance, we revisit our collection of models, plot their average top-1 error vs. validation loss, and notice that error decays exponentially with lower loss (Section 2.4). To describe these observations, we propose a scaling law for downstream error as a function of loss (Section 2.5).
@@ 2.1 Preliminaries
We examine scaling laws that predict the loss $L$ of a model as a function of the compute $C$ in FLOPs used to train the model. If one increases the number of parameters $N$ in a model or the number of tokens $D$ that a model is trained on, compute requirements naturally increase. Hence, $C$ is assumed to be a function of $N,D$. Following Kaplan et al. [49], we use the approximation $C=6ND$, which Hoffmann et al. [43] independently verify. We consider scaling laws,
@@
where $E$ is an irreducible loss and $L^{\prime}$ is the reducible loss. $E$ captures the Bayes error or minimum possible loss achievable on the validation domain. The $L^{\prime}(C)$ term captures what can possibly be learned about the validation domain by training on a source domain. $L^{\prime}(C)$ should go to zero with increased training data and model capacity. $L^{\prime}(C)$ is often assumed to follow a power law: $L^{\prime}(C)=\lambda\cdot C^{-\alpha_{C}}$ (i.a.,  Hestness et al. [41], OpenAI [74]). It is also often helpful to consider a power law in a $\log$-$\log$ plot, where it appears as a line with slope $-\alpha_{C}$ and $y$-intercept $\log{(\lambda)}$.
@@
We define a token multiplier $M=D/N$ as the ratio of training tokens to model parameters. We introduce $M$ for notational convenience as it allows us to consider fixed relationships between $D$ and $N$ even as a model gets bigger (i.e., as $N$ becomes larger).
@@
Hoffmann et al. [43] establish compute-optimal training, where, for any compute budget $H$, the allocation of parameters and tokens that minimize training or validation loss is given by,
@@
To solve for the optimal $N^{*},D^{*}$, one can sweep $N,D$ for each $H$, retaining the best configurations. Hoffmann et al. [43] find that as $H$ increases, $N^{*}$ and $D^{*}$ scale roughly evenly. Assuming equal scaling, there is a fixed compute-optimal token multiplier $M^{*}=D^{*}/N^{*}$ per training distribution.
@@
We define over-training as the practice of allocating compute sub-optimally, so smaller models train on a disproportionately large number of tokens (i.e., $M>M^{*}$). While loss should be higher than in the compute-optimal allocation for a given training budget, the resulting models have fewer parameters and thus are cheaper at inference.
Figure 2: Scaling in the over-trained regime follows consistent power law exponents. We notice parallel lines in the $\log$-$\log$ plots of reducible loss vs. training compute for a range of token multipliers $M$, which give the ratio of training tokens to model parameters and where larger $M$ corresponds to more over-training. For a power law giving reducible loss as a function of compute: $L^{\prime}(C)=\lambda\cdot C^{-\alpha_{C}}$, the scaling exponent $\alpha_{C}$ remains relatively constant resulting in lines with approximately fixed slope. The $y$-intercept, which is determined by the power law coefficient $\lambda$, however, shifts with different values $M$. This suggests that the power law coefficient $\lambda$ is a function of the token multiplier $M$, while the power law exponent $\alpha_{C}$ is not. Hestness et al. [41] report a similar phenomenon of consistent scaling exponents for recurrent networks when modifying architectures instead of token multipliers.
2.2 Observation: Over-trained models follow consistent trends
We begin our scaling investigation by training models with 0.011B to 0.411B parameters for token multipliers $M$ between 20 and 640, where $M=20$ points lie roughly on the compute-optimal frontier, and $M>20$ corresponds to over-training. We defer experimental details to Section 3 to focus on our observations. In Figure 2, we plot loss against compute on a $\log$-$\log$ scale for the models trained on three datasets and evaluated on the C4 eval set. We notice a phenomenon of “parallel lines” when fitting power laws to the reducible loss, which suggests a near-constant scaling exponent even with increased over-training. This indicates that scaling in the over-trained regime may be predictable given training runs closer to compute-optimal.
2.3 Deriving scaling laws for over-trained behavior
In search of an analytic expression for the observations in Figure 2, we turn to the scaling literature. A common functional form for the risk of a model, as proposed in prior work [91, 43] is,
@@ 2.3 Deriving scaling laws for over-trained behavior
Recall from Section 2.1, $N$ is the number of parameters, $D$ the number of training tokens and $E$ the irreducible loss. The constants $E,A,\alpha,B,\beta$ are fit from data. By fitting this parametric form, Hoffmann et al. [43] find that scaling exponents $\alpha$ and $\beta$ are close, suggesting that one should scale $N,D$ equally as compute increases. Hence, we assume $\alpha=\beta$. With this assumption, we reparameterize Equation (3) in terms of compute $C=6ND$ and a token multiplier $M=D/N$. We get the following form,
$L(C,M)=E+\left(aM^{\alpha_{C}}+bM^{-\alpha_{C}}\right)C^{-\alpha_{C}},$ (4)
where $\alpha_{C}=\alpha/2$, $a=A(1/6)^{-\alpha_{C}}$, $b=B(1/6)^{-\alpha_{C}}$ gives the relation to Equation (3). For a complete derivation, see Appendix B.
Equation (4) has the following interpretation: (i) The scaling exponent $\alpha_{C}$ is not dependent on $M$. Thus, we always expect lines with the same slope in the $\log$-$\log$ plot—as in Figure 2. (ii) The term $aM^{\alpha_{C}}+bM^{-\alpha_{C}}$ determines the offsets between curves with different token multipliers. Hence, we expect non-overlapping, parallel lines in the $\log$-$\log$ plot for the range of $M$ we consider—also consistent with Figure 2.
@@
2.4 Observation: Loss tracks average top-1 error
Figure 3: Average top-1 error scales as a function of loss. We plot data points trained on three datasets and notice an exponential decay of average top-1 error as C4 eval loss, on the x-axis, decreases. The specific coefficients appear dataset-dependent. We consider on the y-axes (left) the average error over 46 evaluations and (right) the average error on a subset of 17 evaluations where performance can be 10 points above random chance for at least one 0.154B scale model. These observations suggest that average top-1 error should be predictable with reliable loss estimates.
Scaling is typically studied in the context of loss [49, 43, 70], which Schaeffer et al. [98] note is smoother than metrics like accuracy. However, practitioners ultimately care about downstream, in-the-wild task performance. To better connect scaling laws and over-training to task prediction, we revisit the suite of models plotted in Figure 2. In Figure 3, we plot average downstream top-1 errors over evaluations sourced from LLM-Foundry [67] against the C4 eval loss. We defer details of the setup to Section 3 to focus here on a few key observations. The average errors appear to follow exponential decay as loss decreases. Additionally, the particular relationship between loss and error is dataset-dependent. For instance, models trained on C4 result in the lowest C4 eval loss, but this does not translate to downstream gains compared to models trained on RedPajama or RefinedWeb.
2.5 Proposing a scaling law for average top-1 error
@@
where $\epsilon,k,\gamma$ are fit from data. Equation (5) also has an appealing interpretation in terms of model perplexity $PP(L)=\exp{(L})$,
$\textsf{Err}(PP)=\epsilon-k\cdot PP^{-\gamma}.$ (6)
Namely, Err follows a power law in $PP$ with maximum error $\epsilon$, where intuitively $\epsilon$ should be close to the random chance performance.
Equation (5) in conjunction with (4) suggests a two-step method to predict Err as a function of compute and the amount of over-training. For choices of training and validation distributions, (i) fit a scaling law to Equation (4) using $((C,M),L)$ pairs to yield $(C,M)\mapsto L$. (ii) Fit a scaling law to Equation (5) using $(L,\textsf{Err})$ pairs to get $L\mapsto\textsf{Err}$.
3 Experimental setup
Towards testing the analytic predictions in Equations (4), (5), we discuss our experimental setup. We first present language modeling details (Section 3.1). Next, we discuss our strategy for deciding which models to include in our scaling investigation and our procedure for fitting scaling trends (Section 3.2). We then present metrics to validate how well scaling laws predict loss and downstream performance (Section 3.3).

### 14 Jun 2024 v2

@@ Abstract
Scaling laws are useful guides for derisking expensive training runs, as they predict performance of large models using cheaper, small-scale experiments. However, there remain gaps between current scaling studies and how language models are ultimately trained and evaluated. For instance, scaling is usually studied in the compute-optimal training regime (i.e., “Chinchilla optimal” regime). In contrast, models are often over-trained to reduce inference costs. Moreover, scaling laws mostly predict loss on next-token prediction, but models are usually compared on downstream task performance. To address both shortcomings, we create a testbed of 104 models with 0.011B to 6.9B parameters trained with various numbers of tokens on three data distributions. First, we fit scaling laws that extrapolate in both the amount of over-training and the number of model parameters. This enables us to predict the validation loss of a 1.4B parameter, 900B token run (i.e., 32$\times$ over-trained) and a 6.9B parameter, 138B token run (i.e., a compute-optimal run)—each from experiments that take 300$\times$ less compute. Second, we relate the perplexity of a language model to its downstream task performance by proposing a power law. We use this law to predict top-1 error averaged over downstream tasks for the two aforementioned models, using experiments that take 20$\times$ less compute. Our experiments are available at https://github.com/mlfoundations/scaling.
00footnotetext: Equal advising, ordered alphabetically. Correspondence to sy@cs.columbia.edu. 1Columbia University 2Toyota Research Insitute 3UT Austin 4Apple 5University of Washington 6Juelich Supercomputing Center, Research Center Juelich 7LAION 8Allen Institute for AI 9Stanford University 10Tel Aviv University 11TU Munich 12Contextual AI
@@ 1 Introduction
Figure 1: Reliable scaling with over-training and on downstream error prediction. (left) We fit a scaling law for model validation loss, parameterized by (i) a token multiplier $M=N/D$, which is the ratio of training tokens $D$ to parameters $N$ and (ii) the compute $C$ in FLOPs used to train a model, approximated by $C=6ND$. Larger values of $M$ specify more over-training. We are able to extrapolate, in both $N$ and $M$, the validation performance of models requiring more than 300×300$\times$300 × the training compute used to construct the scaling law. (right) We also fit a scaling law to predict average downstream top-1 error as a function of validation loss. We find that fitting scaling laws for downstream error benefits from using more expensive models when compared to fitting for loss prediction. We predict the average error over 17 downstream tasks for models trained with over 20$\times$ the compute. For this figure, we train all models on RedPajama [112].
@@
Training large language models is expensive. Furthermore, training high-quality models requires a complex recipe of algorithmic techniques and training data. To reduce the cost of finding successful training recipes, researchers first evaluate ideas with small experiments and then extrapolate their efficacy to larger model and data regimes via scaling laws. With reliable extrapolation, it is possible to quickly iterate at small scale and still pick the method that will perform best for the final large training run. Indeed, this workflow has become commonplace for training state-of-the-art language models like Chinchilla 70B [45], PaLM 540B [19], GPT-4 [76], and many others.
@@
Despite their importance for model development, published scaling laws differ from the goals of training state-of-the-art models in important ways. For instance, scaling studies usually focus on the compute-optimal training regime (“Chinchilla optimality” [45]), where model and dataset size are set to yield minimum loss for a given compute budget. However, this setting ignores inference costs. As larger models are more expensive at inference, it is now common practice to over-train smaller models [113]. Another potential mismatch is that most scaling laws quantify model performance by perplexity in next-token prediction instead of accuracy on widely used benchmark datasets. However, practitioners usually turn to benchmark performance, not loss, to compare models.
@@
In this paper, we conduct an extensive set of experiments to address both scaling in the over-trained regime and benchmark performance prediction.
@@
Motivated by the practice of training beyond compute-optimality, we first investigate whether scaling follows reliable trends in the over-trained regime. We notice, as implied by Hoffmann et al. [45], for a set of models of different sizes trained with a constant ratio of tokens to parameters, models’ reducible loss $L^{\prime}$ [43, 45] follows a power law ($L^{\prime}=\lambda\cdot C^{-\eta}$) in the amount of training compute $C$. We find that as one increases the ratio of tokens to parameters, corresponding to more over-training, the scaling exponent $\eta$ remains about the same, while the scalar $\lambda$ changes. We explain our observations by reparameterizing existing scaling laws in relation to the amount of over-training.
@@
To establish empirically that scaling extrapolates in the over-trained regime, we further experiment with a testbed of 104 models, trained from scratch on three different datasets: C4 [88, 27], RedPajama [112], and RefinedWeb [82]. We find that scaling laws fit to small models can accurately predict the performance of larger models that undergo more over-training. Figure 1 (left) illustrates our main over-training result, where we invest $192.4e$ 19 FLOPs to extrapolate the C4 validation performance of a 1.4B parameter model trained on 900B tokens, which requires 300×300$\times$300 × more compute to train.
@@
In addition to over-training, we also investigate if scaling laws can predict the performance of a model on downstream tasks. We establish a power law relationship between language modeling perplexity and the average top-1 error on a suite of downstream tasks. While it can be difficult to predict the error on individual tasks, we find it possible to predict aggregate performance from a model’s perplexity among models trained on the same training data. Figure 1 (right) presents our main downstream error prediction result, where we invest $202.7e$ 20 FLOPs to predict the average top-1 error over a set of downstream tasks to within 1 percentage point for a 6.9B compute-optimal model, which requires 20×20$\times$20 × more compute to train.
@@
Our results suggest that the proposed scaling laws are promising to derisk (i) the effects of over-training models and (ii) the downstream performance of scaling up training recipes. To facilitate further research on reliable scaling, we provide all results of our experiments at https://github.com/mlfoundations/scaling.
2 Developing scaling laws for over-training and downstream tasks
In this section, we develop scaling laws to predict over-trained and downstream performance. First, we provide key definitions (Section 2.1). We next present a scaling law for over-training drawing on empirical observation and prior work (Section 2.2). To connect loss scaling and downstream error prediction, we observe that average top-1 error decreases exponentially as a function of validation loss, which we formalize as a novel scaling law (Section 2.3). In later sections, we build an experimental setup (Section 3) to quantify the extent to which our scaling laws extrapolate reliably (Section 4).
@@ 2.1 Preliminaries
Typically, scaling laws predict model loss $L$ as a function of the compute $C$ in FLOPs used for training. If one increases the number of parameters $N$ in a model or the number of tokens $D$ that a model is trained on, compute requirements naturally increase. Hence, we assume $C$ is a function of $N,D$. Following Kaplan et al. [51], we use the approximation $C=6ND$, which Hoffmann et al. [45] independently verify. We consider,
@@
where $E$ is an irreducible loss and $L^{\prime}$ is the reducible loss. $E$ captures the Bayes error or minimum possible loss achievable on the validation domain. The $L^{\prime}(C)$ term captures what can possibly be learned about the validation domain by training on a source domain. $L^{\prime}(C)$ should approach zero with increased training data and model capacity. $L^{\prime}(C)$ is often assumed to follow a power law: $L^{\prime}(C)=\lambda\cdot C^{-\eta}$ (i.a.,  Hestness et al. [43], OpenAI [76]). It is also often helpful to consider a power law in a $\log$-$\log$ plot, where it appears as a line with slope $-\eta$ and $y$-intercept $\log{(\lambda)}$.
@@
We define a token multiplier $M=D/N$ as the ratio of training tokens to model parameters for notational convenience. $M$ allows us to consider fixed relationships between $D$ and $N$ even as a model gets bigger (i.e., as $N$ becomes larger).
@@
Hoffmann et al. [45] establish compute-optimal training, where, for any compute budget $H$, the allocation of parameters and tokens is given by,
@@
To solve for the optimal $N^{*},D^{*}$, one can sweep $N,D$ for each compute budget, retaining the best configurations. Hoffmann et al. [45] find that as the compute budget increases, $N^{*}$ and $D^{*}$ scale roughly evenly. Assuming equal scaling, there is a fixed compute-optimal token multiplier $M^{*}=D^{*}/N^{*}$ per training distribution.
@@
We define over-training as the practice of allocating compute sub-optimally, so smaller models train on a disproportionately large number of tokens (i.e., $M>M^{*}$). While loss should be higher than in the compute-optimal allocation for a given training budget, the resulting models have fewer parameters and thus incur less inference cost.
Figure 2: Scaling in the over-trained regime follows consistent power law exponents. We notice parallel lines in the $\log$-$\log$ plots of reducible loss vs. training compute for a range of token multipliers $M$, which give the ratio of training tokens to model parameters. Larger $M$ corresponds to more over-training. For a power law giving reducible loss as a function of compute: $L^{\prime}(C)=\lambda\cdot C^{-\eta}$, the exponent $\eta$ remains relatively constant resulting in lines with approximately fixed slope (Figure 17). The scalar $\lambda$ that determines the $y$-intercept, however, shifts with different token multipliers. This suggests $\lambda$ is a function of the token multiplier, while $\eta$ is not.
2.2 Scaling laws for over-training
To propose a scaling law for over-trained models, we first turn to empirical observation. We train four model configurations with parameter counts between 0.011B and 0.411B for token multipliers $M$ between 20 and 640, where $M=20$ points lie roughly on the compute-optimal frontier, and larger $M$ corresponds to more over-training. We defer experimental details to Section 3 to focus on our observations first. In Figure 2, we show loss against compute in a $\log$-$\log$ plot for the models trained on three datasets and evaluated on the C4 eval set. We notice parallel lines when fitting power laws to the reducible loss, which suggests a near-constant scaling exponent even with increased over-training. This indicates that scaling behavior should be describable in the amount of over-training.
In search of an analytic expression for the observations in Figure 2, we consider existing scaling literature. A common functional form for the risk of a model, as proposed in prior work [93, 45] is,
@@ 2.2 Scaling laws for over-training
Recall from Section 2.1, $N$ is the number of parameters and $D$ the number of training tokens. The constants $E,A,\alpha,B,\beta$ are fit from data. By fitting this parametric form, Hoffmann et al. [45] find that scaling exponents $\alpha$ and $\beta$ are roughly equal, suggesting that one should scale $N$ and $D$ equally as compute increases. Hence, we assume $\alpha=\beta$. With this assumption, we reparameterize Equation (3) in terms of compute $C=6ND$ and a token multiplier $M=D/N$. We get,
$L(C,M)=E+\left(aM^{\eta}+bM^{-\eta}\right)C^{-\eta},$ (4)
where $\eta=\alpha/2$, $a=A(1/6)^{-\eta}$, $b=B(1/6)^{-\eta}$ gives the relation to Equation (3). For a complete derivation, see Appendix B.
Equation (4) has the following interpretation: (i) The scaling exponent $\eta$ is not dependent on $M$. Thus, we always expect lines with the same slope in the $\log$-$\log$ plot—as in Figure 2. (ii) The term $aM^{\eta}+bM^{-\eta}$ determines the offsets between curves with different token multipliers. Hence, we expect non-overlapping, parallel lines in the $\log$-$\log$ plot for the range of $M$ we consider—also consistent with Figure 2.
@@
Figure 3: Average top-1 error scales as a function of loss. We plot models trained on three datasets and notice an exponential decay of average top-1 error as C4 eval loss, on the x-axis, decreases. We consider on the y-axes average error on 17 evaluations where performance is at least 10 points above random chance for at least one 0.154B scale model. These observations suggest that average top-1 error should be predictable with reliable loss estimates.
2.3 Scaling laws for downstream error
Scaling is typically studied in the context of loss [51, 45, 72], which Schaeffer et al. [100] note is smoother than metrics like accuracy. However, practitioners often use downstream benchmark accuracy as a proxy for model quality and not loss on perplexity evaluation sets. To better connect scaling laws and over-training to task prediction, we revisit the suite of models plotted in Figure 2. In Figure 3, we plot average downstream top-1 errors over evaluations sourced from LLM-Foundry [69] against the C4 eval loss. We defer details of the setup to Section 3 to focus here on a key observation: average error appears to follow exponential decay as loss decreases.
@@ 2.3 Scaling laws for downstream error
where $\epsilon,k,\gamma$ are fit from data. Equation (5) also has an interpretation in terms of model perplexity $\textsf{PP}(L)=\exp{(L})$,
$\textsf{Err}(\textsf{PP})=\epsilon-k\cdot\textsf{PP}^{-\gamma}.$ (6)
Namely, Err follows a power law in PP that is bounded from above by $\epsilon$ signifying arbitrarily high error and from below by $\epsilon-k\cdot\exp(-\gamma E)$, where $E$ is the Bayes error from Equation (4).
Equation (5) in conjunction with Equation (4) suggests a three-step method to predict Err as a function of compute and the amount of over-training. For choices of training and validation distributions, (i) fit a scaling law to Equation (4) using triplets of compute $C$, token multiplier $M$, and measured loss $L$ on a validation set to yield $(C,M)\mapsto L$. (ii) Fit a scaling law to Equation (5) using pairs of loss $L$ and downstream error Err for models to get $L\mapsto\textsf{Err}$. (iii) Chain predictions to get $(C,M)\mapsto\textsf{Err}$.
3 Constructing a scaling testbed
Figure 4: Search, filter, fit: A recipe for selecting configurations for scaling. (left) To generate the final configurations presented in Table 3, we run a 435 model grid search over model width, hidden dimension, number of attention heads, batch size, and warmup steps. All models are trained near compute-optimally. (center) We plot the efficient frontier of models, which appear to follow a trend, excluding models from $5.2\times 10^{16}$ to $5.2\times 10^{17}$, which fall below the trend. (right) We fit a power law with irreducible error to the remaining configurations, picking four configurations that closely track the full model suite (“Selected models”). These models extrapolate the performance of 1.4B, 6.9B target models. Shaded regions represent bootstrap 95% confidence intervals.
In this section, we discuss our experimental setup to test the predictions suggested by Equations (4) and (5). We first present our general language modeling setup (Section 3.1). Next, we discuss our strategy for determining model configurations for our scaling investigation (Section 3.2) and fitting scaling laws (Section 3.3). We then present metrics to validate how well scaling laws predict loss and downstream performance (Section 3.4).